### v2

Abstract
While automated chemical tools excel at specific tasks, they have struggled to capture the strategic thinking that characterizes expert chemical reasoning. Here we demonstrate that large language models (LLMs) can serve as powerful tools enabling chemical analysis. When integrated with traditional search algorithms, they enable a new approach to computer-aided synthesis that mirrors human expert thinking. Rather than using LLMs to directly manipulate chemical structures, we leverage their ability to evaluate chemical strategies and guide search algorithms toward chemically meaningful solutions. We demonstrate this paradigm through two fundamental challenges: strategy-aware retrosynthetic planning and mechanism elucidation. In retrosynthetic planning, our system allows chemists to specify desired synthetic strategies in natural language – from protecting group strategies to global feasibility assessment – and uses traditional or LLM-guided Monte Carlo Tree Search to find routes that satisfy these constraints. In mechanism elucidation, LLMs guide the search for plausible reaction mechanisms by combining chemical principles with systematic exploration. This approach shows strong performance across diverse chemical tasks, with newer and larger models demonstrating increasingly sophisticated chemical reasoning. Our approach establishes a new paradigm for computer-aided chemistry that combines the strategic understanding of LLMs with the precision of traditional chemical tools, opening possibilities for more intuitive and powerful chemical automation systems.
keywords:
chemical reasoning, large language models, synthesis planning, reaction mechanisms
1 Main
The automation of chemical reasoning has been a long-standing goal in many areas of chemistry, promising to accelerate drug discovery, retrosynthesis[1, 2, 3] and our understanding of chemical reactivity [4]. Traditional machine learning and computational approaches have focused on specialized algorithms for specific tasks – predicting properties [5], planning syntheses [1, 6], or proposing reaction mechanisms [7, 8, 9, 10, 11, 12, 13, 14]. While successful in narrow domains, these systems lack the flexible reasoning and strategic multi-step thinking that characterize expert chemical problem-solving [15, 16, 17]. A synthetic chemist planning the synthesis of a complex molecule, for instance, must simultaneously consider multiple strategic factors: which rings to form first, when to install sensitive functional groups and protective groups, and how to leverage available starting materials [18]. Additionally, this logical deconstruction of molecules must be supported by mechanistic reasoning which requires extrapolation of chemical principles of elementary steps and reactive intermediates, to newly proposed molecules, usually not observed before [4, 19].
The exponential growth in capabilities of Large Language Models (LLMs) has sparked a revolution across scientific disciplines [20, 21, 22, 3], with applications spanning from automated literature analysis to hypothesis generation [23, 24, 25]. These models have demonstrated an unprecedented ability to understand and reason about chemical concepts - from individual functional groups to complete synthetic pathways [3, 26, 27]. Most remarkably, they exhibit reasoning patterns that mirror human chemical intuition rather than traditional computational approaches [28, 21], showing particular promise in analyzing strategic elements of synthesis such as protecting group patterns and ring construction timing [18]. However, a fundamental limitation persists: while LLMs excel at analyzing chemical concepts and strategies, they struggle to generate valid chemical representations, particularly SMILES strings [29, 30, 31], limiting their direct application in critical chemical tasks.
Here, we present a paradigm shift in how LLMs can advance chemical science: rather than attempting to generate chemical structures directly, we position these models as sophisticated reasoning engines that guide traditional search algorithms toward chemically meaningful solutions. This approach combines LLMs’ ability to understand and evaluate complex chemical strategies with the systematic exploration capabilities of established search methods. Through systematic evaluation, we first demonstrate that LLMs can effectively analyze chemical entities and strategic patterns across multiple scales. We then show how these reasoning capabilities can be exploited in two challenging applications: strategy-aware retrosynthetic planning and mechanism elucidation.
Retrosynthetic planning —the process of systematically breaking down complex target molecules into simpler, commercially available starting materials— represents one of the most crucial and intellectually demanding tasks in organic chemistry [1, 32]. Current computational methods employ sophisticated search algorithms guided by either carefully designed or learned heuristics [33, 34, 35]. While these systems excel at finding routes that end in commercially available materials [36, 37], they often struggle to incorporate strategic considerations commonly used by expert chemists, such as optimal timing for ring construction, protecting group introduction or incompatible transformations [18]. Recent specialized systems have addressed specific constraints like starting material availability [38, 39] and bond preservation [40, 41], but a framework for arbitrary strategy-aware synthesis planning has remained elusive. Our approach allows chemists to specify strategic requirements in natural language, with LLM-guided search identifying synthetic routes that satisfy these complex constraints. We demonstrate these capabilities on a challenging benchmark, along with a case study targeting global feasibility analysis and ranking of synthetic routes from multiple sources, including experimentally validated routes.
Figure 1: LLMs as chemical reasoning engines for synthesis planning and mechanism elucidation. a) Current state of LLMs in chemistry, highlighting strengths in property prediction, multiple choice questions, and agentic workflows, alongside limitations in structure generation tasks. b) LLMs demonstrate sophisticated chemical reasoning capabilities, providing detailed analyses of reaction mechanisms and functional group transformations. c) Analytical capabilities go beyond single reactions, being able to analyze a route in global terms. LLMs are capable of detecting when protecting groups are unnecessarily proposed as part of a synthetic route. d) Our key insight: positioning LLMs as strategic evaluators within chemical search frameworks. Rather than generating structures directly, LLMs guide traditional search algorithms toward chemically meaningful solutions. e) Application to synthesis planning: LLMs assess candidate routes based on guidance queries specifying strategic requirements (e.g., ”break pyrimidine in the early stage”). This yields strategically relevant synthetic pathways with detailed rationales for synthetic choices. f) Application to mechanism elucidation: LLMs guide search through possible reaction mechanisms by evaluating the plausibility of elementary electron-pushing steps. The system efficiently identifies correct mechanistic pathways while providing chemically meaningful justifications. This approach combines the strategic understanding of LLMs with the precision of traditional chemical search algorithms.
Similarly, mechanism elucidation —understanding the step-by-step electron movements that transform reactants into products— is fundamental to both chemical understanding and reaction optimization [4, 19, 42]. The power of mechanistic understanding lies not only in explaining individual reactions but in their potential to generalize patterns of chemical reactivity that can be applied to previously unseen molecules [43]. Existing computational methods can enumerate possible reaction paths [7, 44] but often lack the chemical intuition needed to identify plausible mechanisms. While specialized approaches combining quantum calculations with search algorithms show promise [10], they struggle to scale for complex systems and rely on predefined templates or atom-mapping, limiting their applicability. In our approach, basic electron-pushing steps are evaluated using LLMs’ understanding of chemical principles in the context of a search algorithm, allowing the LLM to guide the search towards reasonable mechanisms, while potentially considering diverse forms of experimental evidence and practical constraints. Our results demonstrate that LLMs can effectively guide search processes and select optimal solutions, while providing chemically meaningful rationales for their decisions. Furthermore, we provide insights into how different models’ capabilities —related to pretraining [45], post-training [46, 47, 28] and inference-time scaling [48, 28]— affect solution quality, establishing crucial practical considerations for deploying such systems. This integration represents a significant step toward computational chemistry systems that can reason strategically about complex synthesis challenges while maintaining the precision of traditional computational tools.
2 Results
The approach presented here leverages LLMs as chemical reasoning engines that guide traditional search algorithms through complex chemical spaces. Here, a rule-based process is in charge of generating and proposing intermediate states, while the LLM serves as a judge to select among possible solution paths. The following results and our analyses show that current LLMs are capable of detailed analysis of relevant chemical objects like molecules, reactions, and reactive intermediates. Furthermore, we demonstrate how these capabilities can be leveraged in novel and useful manners for scientific discovery in synthetic chemistry through two challenging applications: prompt-guided and strategy-aware retrosynthetic planning, where natural language queries guide the search for synthetic routes with specific properties, and mechanism elucidation, where the goal is to identify plausible reaction mechanisms by evaluating candidate electron-pushing steps.
2.1 Strategy-aware Synthesis Planning
Retrosynthetic planning represents one of the most challenging tasks in organic chemistry, requiring both deep chemical knowledge and strategic thinking. While computational approaches have successfully automated the search through spaces of reactions, they typically struggle to incorporate the strategic elements that enable synthesis. The complexity of the task arises from multiple challenges across multiple levels and dimensions. At the reaction level, chemists must select transformations that are high-yielding, selective, and technically feasible. At a more global level, long-range considerations make the task almost an art: early synthetic decisions constrain later structural possibilities, protecting groups and functional group interconversions unlock otherwise impossible transformations, and these interconnected choices create intricate decision trees spanning entire synthetic routes. Expert chemists develop sophisticated heuristics for navigating these strategic elements. However, translating this expertise into computational systems has remained challenging.
Here, we extend traditional computer-assisted retrosynthetic search by incorporating natural language specifications of desired synthetic strategies. Given a molecular target and a description of desired route characteristics (e.g. ”construct the pyrimidine ring in early stages”), the system must identify synthetic pathways that satisfy these strategic constraints, see Figure 1e. Such a task requires not only correct understanding of molecular and reaction representations but also an ability to connect chemical theory with practical experimental considerations and modern synthetic methods, depending on the query.
Considering the robust analytical capabilities demonstrated by LLMs (see Figure 1a-c, we developed Synthegy: a framework that combines these analytical capabilities with traditional synthesis planning software (Methods 4.2). For evaluation, we created a benchmark consisting of pairs of molecular targets and prompts, along with scoring scripts that assess route-to-prompt alignment in a tailored manner (Methods 4.4). These prompts specify features of the target synthetic route, and range from simple reaction preferences to complex strategic requirements, allowing us to assess the chemical reasoning capabilities of different LLMs and their ability to effectively guide synthesis planning, see SI-LABEL:si:steering-prompts.
Results in Figure 2a show that current commercial LLMs can already perform advanced reasoning about synthetic routes, successfully evaluating both specific reactions and global strategic features. Large, state-of-the-art models like Gemini-2.5-pro achieve the highest scores. This model achieves such performance by systematically analyzing each reaction in the synthetic sequence, while keeping track of the overall synthetic context, then correlating this with the given prompt (Figure 2b). Our results additionally show that the newer generation of models bring substantial improvements over older generations, with the newest Gemini-2.5-pro performing much better on the harder tasks (Target 4), where no other model could succeed. Figure 2c further highlights the rapid progress that is being achieved thanks to advancements in LLM capability. The plot shows that the kind of intelligence required to tackle the challenge of strategy-aware synthesis planning was entirely missing until around June 2024 with the release of Claude-3.5-Sonnet, which achieves an important improvement in capability relative to any other model at the time. More recently, gemini-2.5-pro established a new state of the art by improving over 50% relative to the next best model. Notably, an open model like DeepSeek-r1 achieves high performances as well, positioning it as a strong open and local alternative to closed providers.
Figure 2: Performance of the system for strategy-aware synthesis planning presented in this work. a Performance of multiple LLMs of different sizes and providers across all the tasks in the benchmark. The tasks are grouped by synthetic target, and each column specifies a prompt as specified in SI-LABEL:si:steering-prompts. The y axis displays the correlation between LLM-produced scores and those computed as specified in the benchmark. Each data point represents a separate execution of the benchmark, with the number of repetitions (n) varying across LLMs based on execution time and resource constraints as specified in the legend. Error bars denote confidence intervals (95%). b Example of an LLM’s analysis of a synthetic route, where it provides a justification why a specific route received a high score. The example illustrates that the LLM analyses each reaction (exemplified with step 4), and then provides an overall analysis where it highlights the alignment with the user’s query. c Average performance of multiple LLMs over time. Shown are only models that are the state of the art for their time, or close to it. The plot illustrates the rapidly evolving advancements in capability. d Illustrates the task of strategy-aware synthesis planning: a user specifies a target molecule along with a query in natural language, which specifies desired features in the route. The proposed solutions are given together with scores that signify their alignment with the query given by the user.
Performance scales strongly with model size. Smaller models’ performance is indistinguishable from random choice even for short routes (see SI-LABEL:app:more-performances), suggesting that the chemical reasoning capabilities required for this task emerge only at larger scales. This suggests either a minimum threshold of model complexity required for meaningful chemical analysis, or a limitation in the multi-task performance of smaller models, suggesting that fine-tuning and other post-processing techniques might be needed to leverage small models on these tasks. Models from other providers (OpenAI [49], DeepSeek [50]) show comparable performance patterns to Claude, with variations potentially attributable to differences in tokenization and prompting strategies rather than fundamental capability gaps.
A key limitation is observed across all models when tackling long synthetic sequences. For the most complex targets (e.g. Target 4, involving routes up to 50 reactions), all models struggle to distinguish and select aligned routes, with the exception of the latest Gemini-2.5-pro [51], which demonstrates unprecedented performance in our benchmark, while producing highly reliable and faithful reasoning traces that enable explainability of the results, see SI-LABEL:app:some-gemini-outputs for examples. Despite these impressive results, some erroneous outputs are still observed. Common failure modes observed across different LLMs include grouping reactions instead of analyzing them individually, which causes overlooking of important fine details, and failing to correctly position individual reactions in the context of a large route. Still, LLMs show overall positive correlations in these extreme cases, showing that some relevant features of routes are still recovered by LLMs, making them useful even for the most complex routes, suggesting their utility as strategic evaluators.
These results demonstrate that current LLMs can effectively assess synthesis planning through analysis and evaluation of arbitrary strategic elements in synthetic routes. In the sections following, we further analyze and validate these findings against real-world chemistry and use-cases, where alternative methodologies can hardly be applied. In particular, we focus on global assessments of feasibility. Under the Synthegy framework, this corresponds to steering towards –or selecting– routes that are overall more feasible, or that are likely to lead to the product in overall high yields.
2.2 Assessing Feasibility of Synthetic Routes
The examples demonstrated in Section 2.1 show the power of Synthegy for selecting synthesis solutions as described in a prompt. The examples shown so far are verifiable by tailored cheminformatics code, thus making their value mostly demonstrative and evaluative, due to the possibility of defining ground truth answers. To test the limits of LLM capability on Synthegy —beyond what current alternative tools allow— we now use the same framework to attempt to select more feasible synthetic routes. Feasibility involves assessing several factors and their interplay: prediction of side-products, hypothesizing and evaluating reaction conditions, feasibility of specific reaction types, functional group incompatibilities, etc. Beyond that, the strategic feasibility of a synthetic route also plays an important role in this regard, as unnecessary reactions can severely hinder overall yields and compromise scalability.
With the same framework as introduced in the previous section, we use the following prompt to find feasible synthetic routes for the targets 1 to 4, see SI-LABEL:si:steering-prompts.
Feasibility prompt Highly feasible synthesis with high overall yields, consider potential side reactions and byproducts. Also ensure no unnecessary reactions are performed.
[…]
Figure 3: Selecting highly feasible routes. a Distribution of feasibility scores as determined by Synthegy, for the four targets in SI-LABEL:si:steering-prompts. Results have been computed for solutions coming from 3 different retrosynthesis engines and one experimentally validated route, for each target. b Example analysis of a feasible synthetic route. The model determines that the two reactions are feasible and the full plan is consistent. c Example of an unfeasible synthetic route. Synthegy correctly identifies 2 key flaws with the presented plan: namely an unfeasible last step, and an inefficient and illogical reaction sequence in preparation for a coupling reaction.
The results in Figure 3a indicate a general bias towards generally infeasible routes in the case of AiZynth, with some small amount of these receiving high scores and thus being deemed feasible. The LLM justifies these low scores with several types of flaws, including ill-conceived protection strategies, reactions with low chances of success, and global strategic flaws such as unnecessary protection cycles and inappropriate ordering of reactions, see SI-LABEL:app:why-low-score for more details. Generally, the commercial tools perform the best as they tend to propose highly feasible solutions in most cases, and in all cases at least one highly feasible route.
Evaluation of the original synthetic routes devised for these targets [54, 55, 56] provide experimental validation of the power of Synthegy at assessing the feasibility of synthetic routes, as these routes were designed by expert chemists and actually executed in a lab, in enough quantities as to allow the evaluation of drug-relevant properties. As shown in Figure 3a, the experimentally realized routes obtain generally high scores, thus providing experimental evidence of the predictive power enabled by Synthegy. Interestingly, results for Target 2 show a general bias towards low scores, even for the experimentally evaluated route, however all tools provide at least one route scoring much higher than the experimentally validated one. Low scores for the experimental route for Target 2 may come from inefficiencies introduced as a result of batch synthesis for screening purposes, where a single route is engineered to synthesize multiple targets with a same scaffold, while varying R groups.
Further inspection of the most feasible solutions reveals that Synthegy enables selection of highly elegant and feasible routes, avoiding protecting group cycles and selecting robust reactions, as shown in Figure 3b. Poorly conceived routes are rapidly discarded by Synthegy (see Figure 3c), citing redundant steps, high potential for side-products and low yields in critical steps, poor regioselectivity, among others, see SI-LABEL:app:why-low-score.
This experiment demonstrates that, given a large and diverse enough space of retrosynthetic solutions for a given target, Synthegy can identify and highlight those that most align with a given query. Such queries can range from simple reaction specifications or relative timing of scaffold formation, to complex feasibility assessments otherwise unfeasible.
2.3 Mechanism Elucidation via LLM-Guided Search
A reaction mechanism is a specification of why and how a given chemical transformation occurs, by means of a set of elementary steps [43]. The power of mechanisms in chemistry lies not only in its explanatory power of a single reaction instance, but also in that the reach of an explanation may extend further than only that reaction; potentially explaining more observed reactions, and also even predicting potential unknown transformations [42].
Figure 4: Mechanisms elucidation a) Requirements and impact of mechanistic elucidation. b) Example of actions in our mechanism framework displayed on example structures, with a post-processing interpretation. c) Example task broken down into moves compliant with our mechanism game framework. d) Model separation performance, averaged across the full mechanism in each case, with n=5 repetitions. e) Comparison of global performance with and without guidance prompts. f) All starting reactants and goal products of our 12 tasks, grouped by category.
We define a set of elementary steps that can be applied to any given molecular set (Methods 4.5 and Figure 4b); these elementary steps can generally be applied to any bond/atom in any molecule, and thus serve as a fundamental basis for formulating mechanisms. The task is then to, given a chemical reaction, find a suitable sequence of such elementary steps that connects the reactants with the products. The search for the best path between reactants and products in chemistry has already been tackled using different ML paradigms, whether using generative models in rule-based environments [8], contrastive learning [13], reinforcement learning in 3D space [57] or flow-matching [14]. Here, we show that LLMs reasoning can be useful for this, even in complex cases. In our approach, an LLM analyzes a partially constructed solution and rates a proposed mechanistic step that continues such solution. For evaluation, we designed a benchmark comprising 12 diverse reactions (Figure 4f) along with their mechanisms (SI-LABEL:si:mechbench-rxns). Performance is measured as the relative score gap between ground truth and alternative moves, averaged over all the steps in the mechanism (Methods 4.7). A perfect scoring model would highly rate correct over incorrect moves, thus this metric correlates with performance in a real search setting by assessing the selectivity of the model at each step.
As shown in Figure 4d, the best model evaluated achieves close to perfect performance on simple reactions like nucleophilic attacks, and poorer performance on more complex tasks like Michael additions on larger molecules and miscellaneous reactions with more complex mechanisms. Despite this drop in performance, the best models can still distinguish good moves over alternative incorrect ones, indicating good adaptability of their chemical knowledge into new situations. Smaller models like gpt-4o-mini generally perform badly, with poor performance even on the simplest of tasks, hinting at similar conclusions as in the previous sections: some complexity threshold might exist after which LLMs become smart enough for being useful at these tasks requiring strong chemical analysis.
Task (# steps) 4o Mini 4o DS-3 Cd 3.5 Cd 3.7 Gm 2.5
Random baseline: $17\%$
1 (4) $50\pm 16$ $55\pm 19$ $75\pm 0$ $90\pm 12$ $\mathbf{100\pm 0}$ $\mathbf{100\pm 0}$
2 (4) $70\pm 10$ $60\pm 12$ $70\pm 10$ $95\pm 10$ $\mathbf{100\pm 0}$ $95\pm 10$
3 (4) $55\pm 10$ $70\pm 10$ $80\pm 10$ $75\pm 0$ $\mathbf{100\pm 0}$ $\mathbf{100\pm 0}$
4 (4) $30\pm 10$ $50\pm 16$ $55\pm 10$ $75\pm 0$ $75\pm 0$ $\mathbf{90\pm 12}$
5 (4) $35\pm 12$ $50\pm 0$ $50\pm 16$ $\mathbf{80\pm 10}$ $\mathbf{80\pm 10}$ $70\pm 19$
6 (8) $25\pm 8$ $52\pm 5$ $70\pm 13$ $80\pm 6$ $\mathbf{85\pm 5}$ $78\pm 39$
7 (12) $20\pm 10$ $48\pm 9$ $20\pm 15$ $75\pm 0$ $\mathbf{80\pm 6}$ $72\pm 37$
8 (8) $5\pm 4$ $18\pm 10$ $28\pm 7$ $\mathbf{68\pm 3}$ $\mathbf{68\pm 10}$ $67\pm 34$
9 (4) $30\pm 10$ $30\pm 10$ $60\pm 25$ $30\pm 10$ $\mathbf{80\pm 10}$ $40\pm 30$
10 (14) $24\pm 10$ $43\pm 6$ $36\pm 12$ $66\pm 3$ $\mathbf{79\pm 5}$ $64\pm 23$
11 (8) $32\pm 6$ $28\pm 9$ $35\pm 5$ $50\pm 8$ $\mathbf{62\pm 8}$ $57\pm 30$
12 (8) $22\pm 9$ $18\pm 6$ $18\pm 6$ $18\pm 6$ $42\pm 6$ $\mathbf{55\pm 19}$
Table 1: Top-1 accuracy on mechanisms elucidation tasks Percentage of ground truth moves classified as strict top-1 when scored with 5 alternatives at each step. Each row displays the results for each of the tasks on the left column. The number of moves involved in each task are specified in parentheses. The reported scores are percentage of correctly scored moves. Scores are reported for LLMs alone, without the help of a guidance prompt defining the mechanism. Each column displays the results for each LLM. The column names used stand for: 4o-mini: GPT-4o-mini, 4o: GPT-4o, DS-3: DeepSeek-3, Cd 3.5: Claude-3.5-Sonnet, Cd 3.7: Claude-3.7-Sonnet, Gm 2.5: Gemini-2.5-pro.
A key advantage of search guided by an LLM is that it allows the specification of any arbitrary amount of information, parameters and instructions, through the LLM’s text interface. Particularly relevant for this task, the input consists of the reaction (reactants and products), but could also contain the reaction conditions (solvent, temperature, concentrations, etc), any available experimental data (such as kinetic studies), among others. In this line of ideas, we experiment with text-guided search, where search is guided by an external text describing the sequence of steps in a hypothetical mechanism. The search here thus functions as a decoder from the input text+reaction, into a proper sequence of elementary steps. The source of this external text can be an expert human, or another LLM which, as is clear from the previous sections, can accurately analyze reactions and describe their mechanisms. Figure 4e shows that adding such information is generally beneficial for performance, boosting the performances of poorly performing LLMs, as is the case of gpt-4o-mini and gpt-4o. Notably, even though Gemini-2.5-Pro’s base performance is already higher than any other competitor, guidance prompts still provide additional and substantial improvements. These results show the potential for including multiple sources of information as input, hinting at the future possibility of automated mechanistic (hypothesis) generation and closed-loop refinement.
3 Discussion
In this work, we demonstrate that LLMs can serve as powerful tools enabling chemical analysis at arbitrary levels of abstraction. We present Synthegy: a framework to exploit these capabilities in scientifically relevant tasks in synthetic chemistry. Our approach combines traditional search algorithms with LLMs’ sophisticated reasoning abilities, positioning the models as expert evaluators that can filter, rerank, or guide search toward solutions that are aligned with chemical principles as well as expert intent, across a wide rang of chemical scales. This integration helps overcome fundamental limitations in existing tools while enabling more intuitive interfaces for complex chemical tasks.
Our results first establish that current LLMs possess remarkable capabilities for detailed chemical reasoning, accurate analysis and evaluation of chemical objects at multiple scales, from functional groups in molecules, to single reactions, to strategic patterns in full synthetic routes. We leverage these capabilities through a framework where candidate solutions are generated by traditionally used computational environments, while LLMs analyze proposed solutions to assess their validity in relation to the specified target solution. Furthermore, these capabilities can be leveraged to directly guide search algorithms towards desirable and chemically-sensible solutions.
We showcase the practical utility of this approach through two challenging applications: strategy-aware synthesis planning and mechanism elucidation. In synthesis planning, our system enables natural language specification of strategic constraints, allowing chemists to select solutions from computer assisted synthesis planning algorithms, towards routes that are more aligned with their intent. Such intent can range from specifications of the reaction sequences used or relative timings of scaffold formation, to full assessments of general feasibility, enabling a novel way of selecting promising and experimentally feasible synthetic routes. We demonstrate this through an experimental study where Synthegy successfully selects the experimentally validated route as highly-feasible, from a non-exhaustive pool of candidate solutions proposed by several standard retrosynthesis tools.
We furthermore evaluate the system on experimentally validated routes for complex targets like Strychnine, and show how the system correctly identifies and evaluates key strategic elements, not only providing key insights into the analyzed solutions but also further validating the practical potential of Synthegy to select solutions with specific properties, even for complex molecular scaffolds and strategic requirements.
Similarly for mechanism elucidation, we show that a similar approach can successfully guide search towards chemically feasible mechanisms, starting only from a set of simple and general elementary steps, making our approach generally applicable to a large space of organic reactions. We note that any arbitrary specification, guidance, or initial condition can be encoded as input to the system, thanks to the flexibility of the LLMs’ text interface. We show the potential of this idea by experimenting with ”guidance prompts” in the input, which resemble intuitions of expert chemists about how the mechanism of a given reaction should look like. These considerations can be extended to include feedback from experiments, that iteratively refine mechanistic hypotheses.
However, some limitations remain. LLMs exhibited a number of failure modes, more prominently eventual misunderstandings of the format of the input data, which drives them to analyze reactions in an inverse sense and thus leading them to deem routes unfeasible. Furthermore, some models display lazyness, which bias models’ responses towards shorter and overly simplistic and optimistic responses. Such behavior can be traced back to pre- and post-training techniques used to train these models. These limitations can be tackled by improved prompting techniques, task-specific fine-tuning, among others. Even with these limitations, we show that the current models are very strong baselines, highlighting the potential for future improvements. Moreover, we show that success in the current implementation is very much dependent on the quality and diversity of synthetic routes generated by the backend retrosynthesis systems. Further research needs to go into the application of these strong capabilities for directly guiding search methods towards solutions that align with prompts and chemical principles, overcoming the inherent limitations of predefined sets of solutions.
Looking forward, our framework opens new possibilities for computer-aided chemistry systems that better align with human chemical intuition while maintaining computational precision. The ability to incorporate diverse forms of chemical knowledge —from initial conditions, to practical constraints, experimental data, or even expert intent— through natural language interfaces could make advanced computational tools more accessible to practicing chemists while enabling new forms of human-AI collaboration in chemical research.
4 Methods
4.1 Large Language Models
Throughout this work we used several LLMs from different providers, all through litellm [58] as it provides a unified interface to multiple provider’s APIs. A temperature of 0.1 has been used across all LLMs. The specific models used throughout this work include: OpenAI models [49] (gpt-4o, gpt-4o-mini, gpt-4-turbo, o3, o4-mini), Anthropic Claude models [59] (claude-3-sonnet-20240229, claude-3-opus-20240229, claude-3-5-sonnet-20241022, claude-3-7-sonnet-20250219, claude-opus-4-20250514, claude-4-sonnet-20250522, claude-2.1) , Meta LLaMA models [60] (llama-3.3-70b-instruct, llama-3.3-8b-instruct, llama-4-scout, llama-4-maverick, llama-3.1-405b-instruct), DeepSeek models [50] (deepseek-r1-0528, deepseek-chat-v3-0324, deepseek-r1-distill variants), Google Gemini models [51] (gemini-2.5-pro-preview, gemini-flash-1.5, gemini-pro-1.5, gemini-2.0-flash-001, gemini-2.5-flash-preview), and Qwen [61] (qwen3-32b).
4.2 Strategy-aware synthesis planning
In this work we present a framework for strategy-aware synthesis planning, which integrates LLMs as evaluators within a traditional retrosynthetic search algorithm. The process begins with a target molecule and a natural language query specifying desired strategic features of the synthetic route. We employed the AiZynthfinder software [16] as the underlying retrosynthesis engine to generate a diverse set of potential synthetic routes for a given target molecule. For each candidate route generated by the underlying retrosynthesis software, we constructed a detailed textual representation of the route, including SMILES strings of all intermediates and reactants, and the sequence of transformations applied at each step. This textual representation, along with the user-provided natural language query, is then passed to the LLM. The LLM is prompted to analyze the synthetic route and evaluate its alignment with the strategic requirements specified in the query. The LLM’s output is a score reflecting the degree to which the route satisfies the user’s query, along with a textual rationale justifying its assessment. This score is then used to rank and filter the candidate routes generated by AiZynthfinder, prioritizing routes that are deemed strategically relevant by the LLM. This process allows us to leverage the systematic exploration capabilities of AiZynth while incorporating the sophisticated chemical reasoning of LLMs to guide the search towards strategically desirable synthetic pathways.
4.3 Feasibility assessment
To evaluate the capability of Synthegy to assess synthetic route feasibility, we collected synthetic routes from multiple sources and compared their LLM-assigned feasibility scores. Routes were obtained from three different retrosynthesis engines alongside experimentally validated routes from the literature.
We accessed synthetic routes through the following platforms:
AiZynth: Synthetic routes were obtained using the AiZynth retrosynthesis planning software [16] with default configuration parameters. The software was run locally using standard installation procedures and default model weights.
Reaxys: Synthetic routes were obtained through the Reaxys API using default parameters as specified by the provider. Access was granted through institutional licensing agreements. The API was queried using target molecule SMILES strings, and the top-ranked routes (by Reaxys scoring) were collected for each target.
Synthia: Routes were generated using Synthia-Lite through their web interface with default parameter settings. No specific parameter customization was performed, utilizing the platform’s standard search algorithms and scoring functions. Routes were collected in SMILES format for subsequent analysis.
Experimental Routes: Literature routes for targets T1-T4 were manually extracted from their respective publications [54, 55, 56] and converted to SMILES format following the same reaction representation standards used for computationally generated routes.
Feasibility Scoring:
All collected routes were evaluated using the Synthegy framework with gemini-2.5-pro as the backend LLM, selected based on its superior performance in our benchmark evaluation (Section 2.1). The feasibility assessment prompt (Section 2.2) was applied consistently across all route sources. Each route was scored independently, with the LLM providing both numerical scores (0-10 scale) and detailed rationales for feasibility assessments.
4.4 Strategy-aware synthesis planning - Benchmark
To evaluate the performance of our strategy-aware synthesis planning framework, we created a benchmark consisting of pairs of molecular targets and prompts. We selected a diverse set of molecular targets that represent various levels of complexity and synthetic challenges. These targets include both well-studied molecules and novel compounds to ensure a comprehensive evaluation. For each molecular target, we crafted prompts that specified desired strategic elements in the synthetic route. These prompts ranged from simple reaction preferences to complex strategic requirements. The specific prompts used are given in the SI LABEL:si:steering-prompts.
As the system returns a set of routes, together with alignment scores, the primary evaluation metric reported in this work is correlation with ground truth scores. The scores are computed using scripts tailored for each target-prompt pair, and work by finding a specifically defined event in a synthetic route. For example, if the prompt requires ”an early ring formation”, then this is translated into a script that finds ring-forming reactions, together with the relative position of this reaction, if any, in the route. A score is thus calculated to account for the happening of the reaction, and it’s relative position, in this case giving a higher score to routes where a ring-forming reaction occurs in the early stages (far from the final product), and progressively lower scores for routes when this happens in later stages. The minimum score is given to routes where the ring-forming reaction doesn’t happen. The full scripts can be found in the github repository.
4.5 Mechanism elementary steps
In order to create an exhaustive and computationally accessible state/action space, mechanisms have been broken down into their most elementary components. Rooted in the arrow-pushing formalism, the possible actions at each state (corresponding to a set of molecules) consist of two fundamental types: ionisation and attack moves. An ionisation move is defined as any bond decreasing its bond order by one, ionising at the same time on one of its terminal atoms. An attack move, on the other hand, is defined as any atom with a lone pair of electrons attacking any atom with an empty orbital, increasing the bond order between them by one, and charging them correspondingly. In both cases, a bond order of 0 corresponds to no bond existing between the two atoms. As minimalistic as this set of moves might seem, it has proven to be quite practical as a systematic way to enumerate and also to translate the majority of non-radical chemistry. However, a limit has been found regarding concerted moves. Even though such cases could often be detected and resolved via a post-processing process, cases like SN1 vs. SN2 (which explicitly require one to know if the electron moves are concerted or not) could not be resolved from the move sequence itself.
4.6 Mechanism elucidation - Benchmark
The mechanism elucidation benchmark was designed to include diverse cases of chemical reactions. It includes several reaction types as well as different scales of molecular weight of the products, assessing different levels of understanding of chemical principles, molecular and reaction representations, and ability to understand the rules proposed by the environment. In Figure 4c and SI-LABEL:si:mechbench-rxns, it is shown how a mechanism is broken down into ionizations and attacks only.
4.7 Mechanism elucidation - Metrics
For each elementary step of the 12 reactions, we computed five alternatives to the ground truth that are legal in our framework. At each step $n$ of a certain mechanism $m$, the LLM is prompted with the mechanistic rules defined here, along with the reactants and products of the reaction, as well as the history of ground-truth moves up to step $n-1$. The prompt also includes a possible next step — one of the 6 in the benchmark — and instructs the LLM to score such move in a scale from 0 to 10 according to alignment with its knowledge of chemical principles and the given reaction. The results presented here correspond to 5 independent runs on each step for reproducibility. The main metric reported in Figure 4d, is the score delta between the ground-truth option, and the mean score for the 5 alternative options, averaged across all steps for a given task. Table 1 reports the percentage of moves for which the ground truth option was given as a strict top-1 (score strictly higher than any of the other possibilities).
Acknowledgments
A.M.B. and P.S. acknowledge support from the NCCR Catalysis (grant number 225147), a National Centre of Competence in Research funded by the Swiss National Science Foundation. T.A.N. acknowleges support from Intel and Merck KGaA via the AWASES programme. D.A. and Z.J. acknowledge support by the Swiss National Science Foundation (SNSF) (grant number 214915). All the authors thank Merck MPS/Synthia and Reaxys for access to their retrosynthetic software and outputs for testing and evaluating the feasibility score.
Data and Code availability
All benchmarks and datasets used, along with the code have been released at https://github.com/schwallergroup/steer. Access to the proprietary LLMs GPT-4, Claude, and DeepSeek can be obtained through each provider’s API. All data generated in this work has been released at https://figshare.com/s/bcf0d2c747e50b7f48d7.